### v1

@@ 7.2.2 Hamilton-Jacobi reachability function
HJ reachability analysis computes the backward reachable set of a constrained system, which is its infeasible region, and safeguards controllers from entering this set. Seo et al. (Seo et al., 2019) use the reachable set obtained by HJ reachability analysis to safeguard receding horizon planning against unknown bounded disturbances. They approximate the reachable set using ellipsoidal parameterization and plan a robust trajectory that avoids risky regions under disturbance. The exact computation of HJ reachability function requires solving an HJ partial differential equation (PDE) on a grid discretization of state space, resulting in an exponential computational complexity with respect to system dimension (Bansal et al., 2017). Many efforts have been made to reduce the computational burden of HJ reachability. Rubies et al. (Rubies-Royo et al., 2019) approximates the optimal controller of HJ reachability problem in control-affine systems as a sequence of simple binary classifiers, thus avoiding storing a representation of HJ reachability function. Herbert et al. (Herbert et al., 2021) propose several techniques, including decomposition, warm-starting, and adaptive grids, to speed up the computation of HJ reachability function. Their methods can update safe sets by one or more orders of magnitude faster than prior work.
@@
Other works further accelerate computation by approximating HJ reachability function with neural networks. Fisac et al. (Fisac et al., 2019) introduce a time-discounted modification of HJ reachability function, which induces a contraction mapping and enables solving it by a fixed point iteration method. Their obtained reachability function approximates the maximum safe set and the safest policy. Based on this work, Yu et al. (Yu et al., 2022a) further consider policy performance optimization in safe RL. They use HJ reachability function to construct virtual-time constraints and solve for the optimal safe policy with the Lagrange method. Bansal et al. (Bansal and Tomlin, 2021) develop a neural PDE solver for high-dimensional reachability problems. The computational requirements of their method do not scale directly with the state dimension but rather with the complexity of the underlying reachable tube.
8 Example of Emergency Braking Control
In this section, we illustrate the proposed concepts, including feasible regions, containment relationships between them, and feasibility function with an emergency braking control problem, where a vehicle is supposed to avoid crash with minimum effort. Fig. 6 gives a schematic of this task.
Figure 6: Emergency braking control scenario.
@@ 8 Example of Emergency Braking Control
The vehicle has a two-dimensional state $x_{t}=[d_{t},v_{t}]^{\top}$ and a one-dimensional action $u_{t}=a_{t}$, where $d$ is the distance to the static obstacle, $v$ is the longitudinal velocity and $a$ is the longitudinal acceleration. It follows a simplified longitudinal dynamics:
$\begin{bmatrix}d_{t+1}\\ v_{t+1}\end{bmatrix}=\begin{bmatrix}1&-\Delta t\\ 0&1\end{bmatrix}\begin{bmatrix}d_{t}\\ v_{t}\end{bmatrix}+\begin{bmatrix}0\\ \Delta t\end{bmatrix}a_{t},$ (13)
@@
where $\Delta t$ is the time step size, assigned with $0.1\mathrm{s}$. We assume a maximum braking deceleration $a_{\mathrm{brk}}=-10\mathrm{m/s^{2}}$ so that the action is bounded in $[a_{\mathrm{brk}},0]$. The reward function (or oppositely in control terminology, the utility function) is
@@
$J=\sum_{i=0}^{\infty}\gamma^{i}r(x_{t+i},u_{t+i})=-\sum_{i=0}^{\infty}\gamma^{i}a_{t+i}^{2}.$ (14)
The safety constraint requires that the vehicle-to-obstacle distance is above zero at all times, requiring collision-free control:
$h(x_{t+i})=-d_{t+i}\leq 0,i=0,1,2,\dots,\infty.$ (15)
The objective (14) follows a general RL setting. To show the applicability of our proposed theoretical framework to both MPC and RL, we also involve an MPC controller in the experiments. The objective of the MPC controller is limited to a finite horizon with length $N=10$ and has no discounting, i.e., $\gamma=1$.
In the following analysis, we build four types of virtual-time constraints: (1) pointwise constraint, (2) CBF constraint, (3) safety index constraint, and (4) HJ reachability constraint. Constraint (1) is a non-feasibility-function-based virtual-time constraint, constraints (2) and (3) are CIS-based, and constraint (4) is CA-based.
@@
$B(x_{i|t})=-d_{i|t}+kv_{i|t}^{2},$ (17)
where $k$ is a tunable parameter. This form is obtained by considering laws of kinematics. Specifically, when braking with a constant acceleration, the distance-to-go is a quadratic function of the current velocity. This form can also be understood from the perspective of HOCBF. If directly choosing the constraint function $h$ as the CBF, its relative degree is two and we should use an HOCBF that considers the first order time derivative of $h$, which is exactly what (17) does. The corresponding CBF constraint applies to the first two steps in virtual-time domain:
$B(x_{0|t})\leq 0,$ (18)
@@
where $\alpha$ is a constant assigned with $0.1$. Strictly speaking, the choice of $\alpha$ is not arbitrary when the action is constrained. It should be chosen such that property 2) in Definition 6.3 is satisfied. Otherwise, $B$ may not be a valid CBF and its zero-sublevel set may not be control invariant. Here, we choose the value of $\alpha$ empirically without verifying property 2).
Safety index follows the form used by Zhao et al. (Zhao et al., 2021):
$\phi(x_{i|t})=\sigma+d_{\mathrm{min}}^{n}-d_{i|t}^{n}+kv_{i|t},$ (19)
where $d_{\mathrm{min}}$ is the minimum allowable distance to obstacle and $\sigma,n,k$ are tunable parameters. In our experiments, we fix $d_{\mathrm{min}}=0,\sigma=0.12$ and adjust the values of $n,k$ to see their effects. $\sigma$ is set slightly larger than zero to avoid small amounts of constraint violation caused by numerical issues in optimization. The safety index constraint applies to the first two steps in virtual-time domain:
$\phi(x_{0|t})\leq 0,$ (20)
@@
where $\eta$ is assigned with 00. Zhao et al. (Zhao et al., 2021) point out that for collision-avoidance tasks with action limits, the following design rule ensures forward invariance of the safe set:
$\frac{n(\sigma+d_{\mathrm{min}}^{n}+kv_{\mathrm{max}})^{\frac{n-1}{n}}}{k}\leq-\frac{a_{\mathrm{brk}}}{v_{\mathrm{max}}},$ (21)
@@
HJ reachability function must be defined with a policy. Here, we consider a safety-oriented policy that takes the maximum braking deceleration at every time step. Its corresponding HJ reachability function is the negative minimum future vehicle-to-obstacle distance starting from the current state:
$F(x_{i|t})=-d_{i|t}-\frac{v_{i|t}^{2}}{2a_{\mathrm{brk}}}.$ (22)
@@
$F(x_{i|t})\leq 0,i=0,1.$ (23)
Note that the CBF (17) is very similar in form to the HJ reachability function (22). In fact, they are equivalent when taking $k=-1/2\cdot a_{\mathrm{brk}}$ in (17). However, the equivalence of the two feasibility functions does not imply the equivalence of their virtual-time constraints. Specifically, the CBF constraint is more restrictive than HJ reachability because of its second constraint for set invariance. This will be clearly shown in the following experiment results.
We solve MPC and RL controllers under these four virtual-time constraints, resulting in eight combinations. For MPC, we use IPOPT to solve the constrained OCPs and compute the optimal actions online. For RL, we adopt the approximate dynamics programming (Li, 2023) framework and use a recently proposed region-wise policy update method called feasible policy improvement (Yang et al., 2023b) to deal with safety constraints. This method solely minimizes constraint violations outside the current feasible region and maximizes rewards under constraints with the interior point method inside the feasible region. It is proved that (Yang et al., 2023b) this method ensures monotonic expansion and convergence of the feasible region.
@@
Figure 7 shows state trajectories of MPC under pointwise constraints with different horizons. Each state trajectory is represented with a solid point followed by a series of hollow points, indicating the initial and successive states, respectively. The hollow point at the end of a trajectory is substituted with a cross if that state violates the real-time constraint indicated by the grey region. The red dashed line is the boundary of the maximum EFR computed analytically by considering the most cautious policy which always brakes with the maximum deceleration. When the constraint horizon $n$ is short, although starting from a state inside the maximum EFR, some trajectories still end up crashing due to the short-sighted control policy under pointwise constraints. This means that the EFR of the optimal policy under pointwise constraints is only a subset of the maximum EFR. As the horizon becomes longer, more states become endlessly feasible under the optimal policy, which coincides with the fact that the maximum EFR actually corresponds to a pointwise constraint with $n=\infty$, as stated in Theorem 4.1.
@@
Figure 8 shows feasible regions of MPC under pointwise constraints with different horizons. Since MPC can be viewed as an optimal policy, its EFR equals that of the optimal policy, and its IFR equals that of the constrained OCP (unrelated to any policy), i.e., $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$. The feasible regions are obtained by running an episode from every state to test its feasibility. Results suggest that short-sighted pointwise constraints fail to recognize the inevitable collision at a distance and, hence, over-optimistically treat some states outside the maximum EFR as initially feasible. At the same time, pointwise constraints are weaker than real-time constraints, so the policy may take too aggressive actions early in the episode, making the EFR of the optimal policy smaller than the maximum EFR. The results also show that as the constraint horizon becomes longer, the EFR of the optimal policy enlarges while the IFR shrinks. When the horizon is long enough, e.g., $n=10$, both regions are almost the same as the maximum EFR. This again supports that with infinite-horizon pointwise constraints, which essentially equal real-time constraints, IFR becomes EFR, and they both equal the maximum EFR.
@@
Figure 7: State trajectories of MPC under pointwise constraints with different horizons. “PW” stands for “pointwise”.
@@
Figure 8: Feasible regions of MPC under pointwise constraints with different horizons.
@@
Figure 9 and 10 show state trajectories and feasible regions of RL under the same pointwise constraint at different iterations. At iteration 10, the feasible regions of the policy are small, and states with high velocities break out of the maximum EFR, resulting in constraint violations. Note that the EFR is smaller than the IFR because, in some states, constraint violation does not happen in the finite virtual horizon (10 steps) but is inevitable in the long run. Here, this infeasibility phenomenon is mainly due to inadequate policy training. As training proceeds, the feasible regions monotonically expand, and more state trajectories are included in the maximum EFR. At iteration 10000, the feasible regions are close to those of the optimal policy of MPC. These results indicate the effectiveness of feasible policy improvement (Yang et al., 2023b), whose theoretical foundation is the feasibility analysis tools for intermediate non-optimal policies proposed in this paper.
@@
Figure 9: State trajectories of RL under the same pointwise constraint ($n=10$) at different iterations.
@@
Figure 10: Feasible regions of RL under the same pointwise constraint ($n=10$) at different iterations.
@@
Figure 11 and 12 show state trajectories and feasible regions of MPC under CBF constraints with different parameters. We observe that the EFR of the optimal policy is always identical to the IFR, which is a common feature of CIS-based virtual-time constraints and is consistent with the analysis in Section 6.2.1 and Corollary 5.1.1(3). A smaller value of $k$ results in a larger feasible region, rendering more states safe in the long run. When $k=0.05$, the CBF actually equals the HJ reachability function (22), and its zero-sublevel set equals the maximum EFR. Another phenomenon is that as the feasible regions expand, restrictions on trajectories inside the feasible regions reduce. This can be seen by comparing the purple or orange trajectory when $k=0.5$ and $k=0.05$, in which the former decelerates immediately while the latter keeps its velocity for some time before decelerating. In other words, trajectories inside the feasible region become less conservative as $k$ decreases.
@@
Figure 11: State trajectories of MPC under CBF constraints with different parameters.
@@
Figure 12: Feasible regions of MPC under CBF constraints with different parameters.
@@
Figure 13 and 14 show state trajectories and feasible regions of RL under the same CBF constraint at different iterations. The feasible regions monotonically expand during training and are close to the maximum EFR at iteration 10000. The state trajectories are either very conservative or infeasible at the beginning and gradually converge to the optimal behavior, i.e., drive the state to the lower left corner. It’s worth noting that the EFR and IFR are nearly always the same during training (we believe their mismatch at iteration 10 is due to inadequate training), which is very different from the case of the safety index and HJ reachability introduced below. This phenomenon is due to the second constraint in (18), which not only requires the state to stay in the feasible region but also restricts the increasing rate of CBF.
@@
Figure 13: State trajectories of RL under the same CBF constraint ($k=0.05$) at different iterations.
@@
Figure 14: Feasible regions of RL under the same CBF constraint ($k=0.05$) at different iterations.
@@
Figure 15 and 16 show state trajectories and feasible regions of MPC under safety index constraints with different parameters. The feasible regions’ boundary shape accurately reflects the safety index’s function form. When $n=2$, the boundary is a quadratic function with respect to $d$. It becomes a linear function when $n=1$ and a square root function when $n=0.5$. When $n=0.5,k=0.23$, the zero-sublevel set of the safety index equals the maximum EFR. Note that when $n=2,k=5$, the EFR is smaller than the IFR because with these parameters, the design rule (21) is violated, and thus, forward invariance of the safe set is not guaranteed. Comparing Figure 15 with Figure 11, we observe a difference between safety index and CBF: under the safety index constraint, the policy does not decelerate until it reaches the boundary of EFR while under the CBF constraint, the policy starts to decelerate before coming close the boundary. This is because of the difference in their constraint design: the safety index only requires the next state to stay in the EFR, while CBF further restricts its value-increasing rate.
@@
Figure 15: State trajectories of MPC under SI constraints with different parameters. “SI” stands for “safety index”.
@@
Figure 16: Feasible regions of MPC under SI constraints with different parameters.
@@
Figure 17 and 18 show state trajectories and feasible regions of RL under the same safety index constraint at different iterations. The EFR monotonically expands and becomes close to the maximum EFR at iteration 10000, while the state trajectories also gradually converge to those of the optimal policy. We observe that at an early stage of training, the IFR under safety index constraint is already quite large, much larger than the EFR. This is different from the case of CBF constraint shown in Figure 14, where EFRs are close to IFRs. The reason is that the safety index constraint only requires the next state to be in the safe set without restricting its value-increasing rate as CBF does. At iteration 10 in Figure 17, the states with high velocities are initially feasible as long as they do not leave the safe set in one step. Accordingly, states close to the safe set boundary are initially infeasible because their next states will leave the safe set.
@@
Figure 17: State trajectories of RL under the same SI constraint ($n=0.5,k=0.23$) at different iterations.
@@
Figure 18: Feasible regions of RL under the same SI constraint ($n=0.5,k=0.23$) at different iterations.
@@
Figure 19 shows state trajectories and feasible regions of MPC under HJ reachability constraint. Being a CA-based virtual-time constraint, the HJ reachability constraint yields identical IFR and EFR of the optimal policy. In addition, since this constraint is built upon the reachability function of a safety-oriented policy, both its IFR and EFR equal the maximum EFR, which coincides with the analysis in Section 6.3.2. This indicates that CA-based virtual-time constraints result in the maximum EFR as long as the optimal feasibility function is found. Comparing the right figure in Figure 19 and the last figure in Figure 11, we observe that the policy under HJ reachability constraint is less conservative than that under CBF constraint, even though the feasible regions of the two constraints are identical. The reason also lies in the fact that CBF constraint restricts the value-increasing rate while HJ reachability constraint does not.
@@
Figure 19: State trajectories and feasible regions of MPC under HJ reachability constraint. “HJR” stands for “HJ reachability”.

### v2

@@ 7.2.2 Hamilton-Jacobi reachability function
HJ reachability analysis computes the backward reachable set of a constrained system, which is its infeasible region, and safeguards controllers from entering this set. While both HJ reachability function and CVF are CA-type feasibility functions, HJ reachability function can better leverage continuous constraint functions such as distance to obstacles, while CVF are typically used with binary cost (constraint violation) signals. This makes HJ reachability function more suitable for accurately representing feasible regions in safety-critical control tasks with continuous state constraints. seo2019robust use the reachable set obtained by HJ reachability analysis to safeguard receding horizon planning against unknown bounded disturbances. They approximate the reachable set using ellipsoidal parameterization and plan a robust trajectory that avoids risky regions under disturbance. The exact computation of HJ reachability function requires solving an HJ partial differential equation (PDE) on a grid discretization of state space, resulting in an exponential computational complexity with respect to system dimension [bansal2017hamilton]. Many efforts have been made to reduce the computational burden of HJ reachability. rubies2019classification approximates the optimal controller of HJ reachability problem in control-affine systems as a sequence of simple binary classifiers, thus avoiding storing a representation of HJ reachability function. herbert2021scalable propose several techniques, including decomposition, warm-starting, and adaptive grids, to speed up the computation of HJ reachability function. Their methods can update safe sets by one or more orders of magnitude faster than prior work.
@@
Other works further accelerate computation by approximating HJ reachability function with neural networks. fisac2019bridging introduce a time-discounted modification of HJ reachability function, which induces a contraction mapping and enables solving it by a fixed point iteration method. Their obtained reachability function approximates the maximum safe set and the safest policy. Based on this work, yu2022reachability further consider policy performance optimization in safe RL. They use HJ reachability function to construct virtual-time constraints and solve for the optimal safe policy with the Lagrange method. bansal2021deepreach develop a neural PDE solver for high-dimensional reachability problems. The computational requirements of their method do not scale directly with the state dimension but rather with the complexity of the underlying reachable tube. To ensure exploration safety under model uncertainty, yu2023safe propose a distributional reachability certificate (DRC) and its Bellman equation to characterize robust persistently safe states, and build a safe RL framework to resolve constraints required by the DRC and its corresponding shield policy. zheng2024safe use a neural HJ reachability function to identify the maximum feasible region in offline RL. They decouple safety constraint adherence, reward maximization, and offline policy learning into three processes, where the optimal policy is derived in the form of weighted behavior cloning.
Chapter 8 Examples of Classic Control Problems
In this chapter, we illustrate the proposed concepts, including feasible regions, containment relationships between them, and feasibility functions with two classic constrained optimal control problems: emergency braking control and unicycle obstacle avoidance. The example code used for all experiments in this chapter is available at our repository [^1] https://github.com/yangyujie-jack/Feasibility-Tutorial.
8.1 Emergency braking control
Figure 8.1: Emergency braking control scenario.
@@ 8.1 Emergency braking control
In emergency braking control, a vehicle is supposed to avoid crash with minimum effort. Figure 8.1 gives a schematic of this task. The vehicle has a two-dimensional state $x_{t}=[d_{t},v_{t}]^{\top}$ and a one-dimensional action $u_{t}=a_{t}$, where $d$ is the distance to the static obstacle, $v$ is the longitudinal velocity and $a$ is the longitudinal acceleration. It follows a simplified longitudinal dynamics:
$\begin{bmatrix}d_{t+1}\\ v_{t+1}\end{bmatrix}=\begin{bmatrix}1&-\Delta t\\ 0&1\end{bmatrix}\begin{bmatrix}d_{t}\\ v_{t}\end{bmatrix}+\begin{bmatrix}0\\ \Delta t\end{bmatrix}a_{t},$ (8.1)
@@
where $\Delta t$ is the time step size, assigned with $0.1\mathrm{s}$. We assume a maximum braking deceleration $a_{\mathrm{brk}}=-10\mathrm{m/s^{2}}$ so that the action is bounded in $a_{\mathrm{brk}},0]$. The reward function (or oppositely in control terminology, the utility function) is
@@
$J=\sum_{i=0}^{\infty}\gamma^{i}r(x_{t+i},u_{t+i})=-\sum_{i=0}^{\infty}\gamma^{i}a_{t+i}^{2}.$ (8.2)
The safety constraint requires that the vehicle-to-obstacle distance be above zero at all times, requiring collision-free control:
$h(x_{t+i})=-d_{t+i}\leq 0,i=0,1,2,\dots,\infty.$ (8.3)
The objective (8.2) follows a general RL setting. To show the applicability of our proposed theoretical framework to both MPC and RL, we also involve an MPC controller in the experiments. The objective of the MPC controller is limited to a finite horizon with length $N=10$ and has no discounting, i.e., $\gamma=1$. The objective for RL is infinite-horizon with $\gamma=0.99$.
In the following analysis, we build four types of virtual-time constraints: (1) pointwise constraint, (2) CBF constraint, (3) SI constraint, and (4) HJ reachability constraint. Constraint (1) is a non-feasibility-function-based virtual-time constraint, constraints (2) and (3) are CIS-based, and constraint (4) is CA-based.
@@
$B(x_{i|t})=-d_{i|t}+kv_{i|t}^{2},$ (8.5)
where $k$ is a tunable parameter. This form is obtained by considering laws of kinematics. Specifically, when braking with a constant acceleration, the distance-to-go is a quadratic function of the current velocity. This form can also be understood from the perspective of HOCBF. If directly choosing the constraint function $h$ as the CBF, its relative degree is two and we should use an HOCBF that considers the first order time derivative of $h$, which is exactly what (8.5) does. The corresponding CBF constraint applies to the first two steps in virtual-time domain:
$B(x_{0|t})\leq 0,$ (8.6)
@@
where $\alpha$ is a constant assigned with $0.1$. Strictly speaking, the choice of $\alpha$ is not arbitrary when the action is constrained. It should be chosen such that property 2) in Definition 6.2.1 is satisfied. Otherwise, $B$ may not be a valid CBF and its zero-sublevel set may not be control invariant. Here, we choose the value of $\alpha$ empirically without verifying property 2).
SI follows the form used by zhao2021model:
$\phi(x_{i|t})=\sigma+d_{\mathrm{min}}^{n}-d_{i|t}^{n}+kv_{i|t},$ (8.7)
where $d_{\mathrm{min}}$ is the minimum allowable distance to obstacle and $\sigma,n,k$ are tunable parameters. In our experiments, we fix $d_{\mathrm{min}}=0,\sigma=0.12$ and adjust the values of $n,k$ to see their effects. $\sigma$ is set slightly larger than zero to avoid small amounts of constraint violation caused by numerical issues in optimization. The SI constraint applies to the first two steps in virtual-time domain:
$\phi(x_{0|t})\leq 0,$ (8.8)
@@
where $\eta$ is assigned with 0. zhao2021model point out that for collision-avoidance tasks with action limits, the following design rule ensures forward invariance of the safe set:
$\frac{n(\sigma+d_{\mathrm{min}}^{n}+kv_{\mathrm{max}})^{\frac{n-1}{n}}}{k}\leq-\frac{a_{\mathrm{brk}}}{v_{\mathrm{max}}},$ (8.9)
@@
HJ reachability function is intrinsically associated with a policy, which is typically the optimal one derived from dynamic programming, but can also be a pre-specified policy. Here, we consider a safety-oriented policy that takes the maximum braking deceleration at every time step. Its corresponding HJ reachability function is the negative minimum future vehicle-to-obstacle distance starting from the current state:
$F(x_{i|t})=-d_{i|t}-\frac{v_{i|t}^{2}}{2a_{\mathrm{brk}}}.$ (8.10)
@@
$F(x_{i|t})\leq 0,i=0,1.$ (8.11)
Note that the CBF (8.5) is very similar in form to the HJ reachability function (8.10). In fact, they are equivalent when taking $k=-1/2\cdot a_{\mathrm{brk}}$ in (8.5). However, the equivalence of the two feasibility functions does not imply the equivalence of their virtual-time constraints. Specifically, the virtual-time constraint constructed by the CBF is more restrictive than that constructed by the HJ reachability function, since the second constraint in the CBF case requires $B(x_{1|t})\leq(1-\alpha)B(x_{0|t})$ where $(1-\alpha)B(x_{0|t})$ may be a negative number, while that in the HJ reachability case only requires $F(x_{1|t})\leq 0$.
We solve MPC and RL controllers under these four virtual-time constraints, resulting in eight combinations. For MPC, we use IPOPT to solve the constrained OCPs and compute the optimal actions online. For RL, we adopt the approximate dynamic programming [li2023reinforcement] algorithm, which trains a policy network and a value network. Both networks are multilayer perceptrons with two hidden layers of 64 neurons each and ReLU activation functions. The value network is updated by minimizing the mean squared error between its prediction and the target computed by one-step look-ahead. The loss function is
$\mathcal{L}_{V}(\phi)=\mathbb{E}\left[\left(V_{\phi}(x_{t})-\left(r(x_{t},u_{t})+\gamma V_{\phi}(x_{t+1})\right)\right)^{2}\right],$ (8.12)
where $\phi$ are the parameters of the value network. The policy network is updated by feasible policy improvement [yang2023feasible], which splits the state space into feasible region and infeasible region, and applies different update rules in these two regions. Inside the feasible region, we maximize the value function:
$\mathcal{L}_{\pi,\text{in}}(\theta)=-\mathbb{E}\left[\mathbb{I}_{\mathrm{X}_{\mathrm{edls}}^{g}}(x_{t})\cdot(r(x_{t},u_{t})+\gamma V_{\phi}(x_{t+1}))\right],$ (8.13)
where $\theta$ are the parameters of the policy network, and $\mathbb{I}_{\mathrm{X}_{\mathrm{edls}}^{g}}$ is the indicator function of the feasible region. Outside the feasible region, we minimize the feasibility function:
$\mathcal{L}_{\pi,\text{out}}(\theta)=\mathbb{E}\left[(1-\mathbb{I}_{\mathrm{X}_{\mathrm{edls}}^{g}}(x_{t}))\cdot F(x_{t+1})\right],$ (8.14)
where $F$ is the feasibility function used in the virtual-time constraint. The total loss for policy network is
$\mathcal{L}_{\pi}(\theta)=\mathcal{L}_{\pi,\text{in}}(\theta)+\mathcal{L}_{\pi,\text{out}}(\theta).$ (8.15)
It is proved [yang2023feasible] that this method guarantees monotonic expansion of the feasible region, monotonic improvement of the value function, and convergence to the optimal policy. The optimizer for both networks is Adam, and the learning rate is $10^{-4}$. A mini-batch of size 256 is sampled for each update. More implementation details can be found in our source code.
@@
Figure 8.2 shows state trajectories of MPC under pointwise constraints with different horizons. Each state trajectory is represented with a solid point followed by a series of hollow points, indicating the initial and successive states, respectively. The hollow point at the end of a trajectory is substituted with a cross if that state violates the real-time constraint indicated by the grey region. The red dashed line is the boundary of the maximum EFR computed analytically by considering the most cautious policy which always brakes with the maximum deceleration. When the constraint horizon $n$ is short, although starting from a state inside the maximum EFR, some trajectories still end up crashing due to the short-sighted control policy under pointwise constraints. This means that the EFR of the optimal policy under pointwise constraints is only a subset of the maximum EFR. As the horizon becomes longer, more states become endlessly feasible under the optimal policy, which coincides with the fact that the maximum EFR actually corresponds to a pointwise constraint with $n=\infty$, as stated in Theorem 4.2.4.
@@
Figure 8.3 shows feasible regions of MPC under pointwise constraints with different horizons. Since MPC can be viewed as an optimal policy, its EFR equals that of the optimal policy, and its IFR equals that of the constrained OCP (unrelated to any policy), i.e., $\mathrm{X}_{\mathrm{init}}^{g}(\pi^{*})=\mathrm{X}_{\mathrm{init}}^{g}$. The feasible regions are obtained by running an episode from every state to test its feasibility. Results suggest that short-sighted pointwise constraints fail to recognize the inevitable collision at a distance and, hence, over-optimistically treat some states outside the maximum EFR as initially feasible. At the same time, pointwise constraints are weaker than real-time constraints, so the policy may take too aggressive actions early in the episode, making the EFR of the optimal policy smaller than the maximum EFR. The results also show that as the constraint horizon becomes longer, the EFR of the optimal policy enlarges while the IFR shrinks. When the horizon is long enough, e.g., $n=10$, both regions are almost the same as the maximum EFR. This again supports that with infinite-horizon pointwise constraints, which essentially equal real-time constraints, IFR becomes EFR, and they both equal the maximum EFR.
@@
Figure 8.2: State trajectories of MPC under pointwise constraints with different horizons. “PW” stands for “pointwise”.
@@
Figure 8.3: Feasible regions of MPC under pointwise constraints with different horizons.
@@
Figure 8.4 and 8.5 show state trajectories and feasible regions of RL under the same pointwise constraint at different iterations. At iteration 10, the feasible regions of the policy are small, and states with high velocities break out of the maximum EFR, resulting in constraint violations. Note that the EFR is smaller than the IFR because, in some states, constraint violation does not happen in the finite virtual horizon (10 steps) but is inevitable in the long run. Here, this infeasibility phenomenon is mainly due to inadequate policy training. As training proceeds, the feasible regions monotonically expand, and more state trajectories are included in the maximum EFR. At iteration 10000, the feasible regions are close to those of the optimal policy of MPC. These results indicate the effectiveness of feasible policy improvement [yang2023feasible], whose theoretical foundation is the feasibility analysis tools for intermediate non-optimal policies proposed in this paper.
@@
Figure 8.4: State trajectories of RL under the same pointwise constraint ($n=10$) at different iterations.
@@
Figure 8.5: Feasible regions of RL under the same pointwise constraint ($n=10$) at different iterations.
@@
Figure 8.6 and 8.7 show state trajectories and feasible regions of MPC under CBF constraints with different parameters. We observe that the EFR of the optimal policy is always identical to the IFR, which is a common feature of CIS-based virtual-time constraints and is consistent with the analysis in Section 6.2.1 and Corollary 5.1.3(3). A smaller value of $k$ results in a larger feasible region, rendering more states safe in the long run. When $k=0.05$, the CBF actually equals the HJ reachability function (8.10), and its zero-sublevel set equals the maximum EFR. Another phenomenon is that as the feasible regions expand, restrictions on trajectories inside the feasible regions reduce. This can be seen by comparing the purple or orange trajectory when $k=0.5$ and $k=0.05$, in which the former decelerates immediately while the latter keeps its velocity for some time before decelerating. In other words, trajectories inside the feasible region become less conservative as $k$ decreases.
@@
Figure 8.6: State trajectories of MPC under CBF constraints with different parameters.
@@
Figure 8.7: Feasible regions of MPC under CBF constraints with different parameters.
@@
Figure 8.8 and 8.9 show state trajectories and feasible regions of RL under the same CBF constraint at different iterations. The feasible regions monotonically expand during training and are close to the maximum EFR at iteration 10000. The state trajectories are either very conservative or infeasible at the beginning and gradually converge to the optimal behavior, i.e., drive the state to the lower left corner. It’s worth noting that the EFR and IFR are nearly always the same during training (we believe their mismatch at iteration 10 is due to inadequate training), which is very different from the case of the SI and HJ reachability introduced below. This phenomenon is due to the second constraint in (8.6), which not only requires the state to stay in the feasible region but also restricts the increasing rate of CBF.
@@
Figure 8.8: State trajectories of RL under the same CBF constraint ($k=0.05$) at different iterations.
@@
Figure 8.9: Feasible regions of RL under the same CBF constraint ($k=0.05$) at different iterations.
@@
Figure 8.10 and 8.11 show state trajectories and feasible regions of MPC under SI constraints with different parameters. The feasible regions’ boundary shape accurately reflects the SI’s function form. When $n=2$, the boundary is a quadratic function with respect to $d$. It becomes a linear function when $n=1$ and a square root function when $n=0.5$. When $n=0.5,k=0.23$, the zero-sublevel set of the SI equals the maximum EFR. Note that when $n=2,k=5$, the EFR is smaller than the IFR because with these parameters, the design rule (8.9) is violated, and thus, forward invariance of the safe set is not guaranteed. Comparing Figure 8.10 with Figure 8.6, we observe a difference between SI and CBF: under the SI constraint, the policy does not decelerate until it reaches the boundary of EFR while under the CBF constraint, the policy starts to decelerate before coming close the boundary. This is because of the difference in their constraint design: the SI only requires the next state to stay in the EFR, while CBF further restricts its value-increasing rate.
@@
Figure 8.10: State trajectories of MPC under SI constraints with different parameters.
@@
Figure 8.11: Feasible regions of MPC under SI constraints with different parameters.
@@
Figure 8.12 and 8.13 show state trajectories and feasible regions of RL under the same SI constraint at different iterations. The EFR monotonically expands and becomes close to the maximum EFR at iteration 10000, while the state trajectories also gradually converge to those of the optimal policy. We observe that at an early stage of training, the IFR under SI constraint is already quite large, much larger than the EFR. This is different from the case of CBF constraint shown in Figure 8.9, where EFRs are close to IFRs. The reason is that the SI constraint only requires the next state to be in the safe set without restricting its value-increasing rate as CBF does. At iteration 10 in Figure 8.12, the states with high velocities are initially feasible as long as they do not leave the safe set in one step. Accordingly, states close to the safe set boundary are initially infeasible because their next states will leave the safe set.
@@
Figure 8.12: State trajectories of RL under the same SI constraint ($n=0.5,k=0.23$) at different iterations.
@@
Figure 8.13: Feasible regions of RL under the same SI constraint ($n=0.5,k=0.23$) at different iterations.
@@
Figure 8.14 shows state trajectories and feasible regions of MPC under HJ reachability constraint. Being a CA-based virtual-time constraint, the HJ reachability constraint yields identical IFR and EFR of the optimal policy. In addition, since this constraint is built upon the reachability function of a safety-oriented policy, both its IFR and EFR equal the maximum EFR, which coincides with the analysis in Section 6.3.2. This indicates that CA-based virtual-time constraints result in the maximum EFR as long as the optimal feasibility function is found. Comparing the right figure in Figure 8.14 and the last figure in Figure 8.6, we observe that the policy under HJ reachability constraint is less conservative than that under CBF constraint, even though the feasible regions of the two constraints are identical. The reason also lies in the fact that CBF constraint restricts the value-increasing rate while HJ reachability constraint does not.
@@
Figure 8.14: State trajectories and feasible regions of MPC under HJ reachability constraint. “HJR” stands for “HJ reachability”.